Hawk Homes, Inc., makes one type of birdhouse that it sells for $30.40 each. Its variable cost is $14.90 per house, and its fHawk Homes, Inc., makes one type of birdhouse that it sells for $30.40 each. Its variable cost is $14.90 per house, and its fHawk Homes, Inc., makes one type of birdhouse that it sells for $30.40 each. Its variable cost is $14.90 per house, and its fHawk Homes, Inc., makes one type of birdhouse that it sells for $30.40 each. Its variable cost is $14.90 per house, and its fHawk Homes, Inc., makes one type of birdhouse that it sells for $30.40 each. Its variable cost is $14.90 per house, and its f
Answer #1

1) Contribution margin income statement

Total Per unit
Sales 82080 30.40
Variable cost 40230 14.90
Contribution margin 41850 15.50
Fixed cost 13314.50
Operating income 28535.50

2) Total Contribution margin = $13314.50

3) Contribution margin per unit = 30.40-14.90 = 15.50

Contribution margin ratio = 15.50/30.40 = 51%

4) Break even unit = 13314.50/15.5 = 859 Units

Break even revenue = 859*30.40 = 26113.60

5) Target unit sale = (13314.50+22000)/15.50 = 2278 Units